OpenStudy (anonymous):

find the slope of (-4), 2 (and) (1,-3)

OpenStudy (texaschic101):

slope formula is : slope = (y2 - y1) / (x2 - x1) (-4,2)...x1 = -4 and y1 = 2 (1,-3)...x2 = 1 and y2 = -3 now we sub slope = (-3 - 2) / (1 - (-4) slope = -5/(1 + 4) slope = -5/5 slope = -1

OpenStudy (anonymous):

hope u got it